A 10 kg ball traveling 2 m/s collides elastically with a 2 kg ball traveling 4 m/s. After the collision the 2 kg ball rolls away at 6m/s. Find the velocity of the 10 kg ball.

m1v1 + m2v2 before= m1v1 + m2v2 after
1
Expert's answer
2020-11-10T07:02:40-0500
"m_1v_1 -m_2v_2=m_1v_1'+ m_2v_2'\\10(2)-2(4)=10v_1'+2(6)\\v_1'=0\frac{m}{s}"